An electron has the following wave function: ψ(x)=A(e^-x)(1-e^-x) Determine i) the normalization constant A, and ii) the expectation value of x .
1
Expert's answer
2020-04-23T12:58:26-0400

The wave function of an electron

ψ(x) = Ae−x(1−e−x)


i) The normalization condition

1 = "\int^∞_0∣ψ(x)∣"2dx

1 = A2 "\u222b _\n0^\n\u221e\n\u200b\t\n e"-2x(1 - e-x)2dx = A2 / 12

A = 2"\sqrt{3}"


(ii) The expectation value of x

⟨x⟩ = 12"\int_0^∞"xe−2x(1−ex)2dx = 13 / 12 ​= 1.08"\u200b"